OpenStudy (anonymous):

solve: x^2-4x=5 show steps? :) thank you

OpenStudy (anonymous):

It's a quadratic equation, isn't it? :) Factorize or use the formula. :)

OpenStudy (aroub):

x^2-4x-5=0 (x-5)(x+1)=0 x=5 x=-1

OpenStudy (anonymous):

Thanks to both :)

OpenStudy (aroub):

no problem! :D